OpenStudy (anonymous):

What is the slope and y-intercept of this line? y = –3x + 9

OpenStudy (justuu):

What do you think?

OpenStudy (anonymous):

I tried to find the slope and I got 3, but I don't do slope very well :(

OpenStudy (anonymous):

do you know what form this equation is in?

OpenStudy (anonymous):

Is it in slope intercept, standard form, or point slope form?

OpenStudy (anonymous):

Slope intercept

OpenStudy (anonymous):

what is Slope int form? y=mx+b?

OpenStudy (anonymous):

I believe so

OpenStudy (anonymous):

What do you think the m stands for in slope intercept form?

OpenStudy (anonymous):

m, x and b have meanings.

OpenStudy (anonymous):

I don't know, that's mainly what I don't understand. I'm homeschooled, so I don't have a very smart teacher online.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

y=mx+b is my favorite form of an equation.

OpenStudy (anonymous):

b=y-intercept in all y=mx+b equations

OpenStudy (anonymous):

If you ever run into an equation with no b, then the y intercept is 0 (this only applies to Slope int form equations)

OpenStudy (anonymous):

isn't it ax+by=c? or no

OpenStudy (anonymous):

ax=by=c is standard form.

OpenStudy (anonymous):

oh,

OpenStudy (anonymous):

y = –3x + 9 Okay, back to this.

OpenStudy (anonymous):

y=-3x+9 **

OpenStudy (anonymous):